OpenStudy (anonymous):

Bev and mike drove 140 miles before stopping at a rest area. after leaving the rest area they traveled an additional 90 miles at an average speed that was 4 miles per hours faster than their speed before stopping. If the entire trip took 4 hours of driving time, what was their average speed during each part of the trip?

OpenStudy (anonymous):

Solve the following equation for r, the average speed for the first 140 miles:\[\frac{140}{r}+\frac{90}{r+4}= 4\]r = 56mph